OpenStudy (anonymous):

Combine like terms. You may use the properties to help: 16 – 5q – 2q –q

OpenStudy (stamp):

pretty straightforward request. combine the like terms.

OpenStudy (anonymous):

16-8q?

OpenStudy (stamp):

ok

OpenStudy (anonymous):

-8q + 16 = 0 -8q = -16 -8q/-8 = -16/-8 q = 2

OpenStudy (stamp):

@choopers you are doing it wrong

OpenStudy (anonymous):

these were the possible answers 16 – 9q 16 – 7q 16 -10q 16 – 8q

OpenStudy (stamp):

belengee you have the answer you already did it this is a closed question...

OpenStudy (anonymous):

16 – 8q sorry. I thought you were looking for the value of q. :)))

OpenStudy (anonymous):

its all good. i appreciate the help tho.